OpenStudy (anonymous):

4^k=x^2+8x please help solve in terms of x please.

OpenStudy (anonymous):

Okay so, its a quadratic equation because the highest power of x is 2. So lets 'complete the square'. \[x^2+8x=4^k\] So add half the coefficient of x to both side, so the equation is still equal. \[(x^2+8x+4^2)=4^k+4^2\] Now its a perfect square. So therefore: \[(x+4)^2=4^k+16\] Now we just simplify. \[x+4=\pm\sqrt(4k+16)\] And again. \[x=-4\pm4\sqrt(4^k)\]

OpenStudy (anonymous):

thank youuuu!

OpenStudy (anonymous):

oh right i meant "add the square of half the coefficient of x". :-)